OpenStudy (anonymous):

use the elimination method to solve the system of equations. 5u+3v=-17 2u+v=-8 please show work. thanks

hero (hero):

v = (-17-5u)/3 v = -8 - 2u Set v = v (-17-5u)/3 = -8 - 2u -17-5u = 3(-8-2u) -17-5u = -24 -6u u = -7 v = 6

hero (hero):

Any questions?

OpenStudy (anonymous):

No, thanks